A stone is released freely from a 200 m tall building. Calculate the final veocity of the stone just before touching the ground. Ignore the friction of air. (use g = 9.81 ms-1)
1
Expert's answer
2020-03-27T09:11:28-0400

From the conservation of energy:


"0.5mv^2=mgh"

"v^2=2gh\to v^2=2(9.81)(200)"

"v=62.6\frac{m}{s}"